a bowl contains 75 red jelly beans and 60 black jelly beans. if 20 black jelly beans are added to the bowl, determine the number of red jelly beans that must be
added to make the ratio of red to black equal to 3:2
a 0 b 20 c 30 d 45 e 60

Answer by stanbon(57354)

a bowl contains 75 red jelly beans and 60 black jelly beans. if 20 black jelly beans are added to the bowl, determine the number of red jelly beans that must be added to make the ratio of red to black equal to 3:2
----
Let number of red that must be added be "x":
(75+x)/80 = 3/2
75+x = 120
x = 45
============
Cheers,
Stan H.
============

a 0 b 20 c 30 d 45 e 60